C1 C3 C5 C2 C4

C1=4F, C2=4F, C3=2F, C4=4F, C5= 14.3 F.

Calculate the equivalent capacitance between A and B points.
Answer #1

since C1 abd C2 are in parallel the net capacitece

C' = C1 + C2 = 4 + 4 = 8 F

C3 abd C4 are in parallel the net capacitece

C'' = C3 + C4 = 2+ 4 = 6 F

now C' , C'' and C5 are in series

the equivalent capacitance between A and B points is

1/Ceq = 1/C' + 1/C'' +1/ C5

= 1/8+1/6 + 1/14.3

Ceq = 2.76 F